OpenStudy (anonymous):

Is there anyone who is good in Geometry who can guide me through a few questions?

OpenStudy (anonymous):

That depends......I may have to report you.. lol jk:)

OpenStudy (anonymous):

yea...lol

OpenStudy (anonymous):

Ask away....

OpenStudy (anonymous):

4. The area of a regular octagon is 35 cm². What is the area of a regular octagon with sides three times as long? (1 point)

OpenStudy (anonymous):

315 cm² 175 cm² 225 cm² 105 cm²

OpenStudy (nathan917):

I did this question yesterday. 8. The area of an octagon is a²(2+2√2) where a is side length So we know 35 = a²(2+2√2) A = (8a)²(2+2√2) Rearrange the first in terms of a² a² = 35 / (2+2√2) Substitute in the second and the (2+2√2) terms cancel out A = 64 * 35 = 2,240cm²

OpenStudy (anonymous):

The area increases with the square of the side length, so a 3x increase in side length corresponds to a 9x increase in area. The new area is 35x9=315cm2.

OpenStudy (anonymous):

What do you mean you did this question yesterday

OpenStudy (anonymous):

@nathan917

OpenStudy (anonymous):

15 in.² 259.8 in.² 519.6 in.² 52 in.² 5. Find the area of the regular hexagon with a side of 10 in. Give the answer to the nearest tenth. (1 point)

OpenStudy (anonymous):

15 in.² 259.8 in.² 519.6 in.² 52 in.²

OpenStudy (anonymous):

heeloevery1 aap chutiya hai..ask the meaning to any indian :P

OpenStudy (nathan917):

I was just saying That i did it yesterday and you answer are different. 35 cm^2 x 3 x 3 = 35 cm^2 x 9 = 315 cm^2

OpenStudy (anonymous):

@crazycatlady are you still there?

OpenStudy (anonymous):

I'm back. Went to get coffee..Banana nut mmmm (: